# Can xLLMs Understand the Structure of Dialog? Exploring Multilingual Response Generation in Complex Scenarios

### v1

@@ 5.1 Dataset Split and Baseline Models
To benchmark the performance of state-of-the-art large language models (LLMs) on the XMP dataset, we selected three open-source instruction-tuned models as baselines: LLaMA 3.1 8B Instruct, Qwen 2.5 7B Instruct, and LLaMA 3.1 70B Instruct. These models were chosen for their demonstrated capabilities in multilingual dialogue response task. Their inclusion allows us to examine the trade-offs between model size, and response quality across six languages
@@ 5.2 Task Setup
The generation task focuses on producing high-quality responses in complex multi-party dialogue scenarios. Given a conversational context and the current speaker, the model is tasked with generating a coherent, contextually relevant, and linguistically accurate response.
Specially, Let $D={(H,S,R)}^{n}_{i=1}$represent the multilingual dialogue dataset for complex scenarios, where $H$ is the dialogue history composed of multiple “Speaker: Utterance” pairs concatenated sequentially, $S$ is the current speaker, and $R$ is the ground truth response. To guide the model in generating speaker-specific responses, a prompt is dynamically constructed based on the current speaker, such as “What will Speaker-x say next?”. The dialogue history is then concatenated with this prompt to form the input for the model. During training, the model predicts the response $\hat{R}$, and the cross-entropy loss is computed between and the ground truth $L=-\frac{1}{m}{\textstyle\sum_{j=1}^{m}}R_{j}log\hat{R}_{j}$, where $m$ is the number of tokens in $R$.
@@ 5.3 Implementation Details
We utilized LLaMAFactory as the training framework and applied LoRA (Low-Rank Adaptation) for efficient fine-tuning. The training was conducted on NVIDIA A100 GPUs using the AdamW optimizer with a learning rate of $1e^{-5}$, a warm-up of 500 steps, and a batch size of 8. To enhance generalization, we standardized key entities in the dialogue history and replaced real entity data with domain-specific slot placeholders. For inference, we used vLLM, an optimized framework that ensures fast and scalable generation. Detailed training and inference commands, along with parameter settings, are provided in Appendix E.
@@ 6.1 Model Performance
The performance gain of LLaMA-3.1 70B over smaller models (e.g., 8B or Qwen-2.5 7B) is marginal, with improvements typically within 1% across most metrics. This suggests that the emergent abilities often associated with larger models are not prominently reflected in this task. Instead, the slight advantage may stem from the larger models’ capacity to generate grammatically and lexically refined text, rather than demonstrating a better understanding of the task.
Additionally, Japanese yields unexpectedly high scores across all models. Despite its low proportion in pretraining corpora, Japanese outperforms other non-English languages. We hypothesize that this stems from the unique syntactic structure and word composition of Japanese, which relies heavily on fixed sentence patterns and the use of kana. Tokenization using Janome likely also contributes to higher n-gram overlap, boosting surface-level metrics like BLEU and ROUGE. To validate this hypothesis, we included BERTScore as an additional evaluation metric. Unlike n-gram metrics, BERTScore evaluates semantic similarity.
6.2 Semantic Analysis with BERTScore
@@
Table 4 presents BERTScore results for all six languages. English achieves the highest semantic alignment, confirming its strong representation in training data. Japanese, despite its high BLEU and ROUGE scores, shows comparable BERTScore results to other non-English languages, indicating that its apparent advantage in traditional metrics stems from tokenization effects rather than true semantic superiority.
For other languages, the results indicate that Japanese achieves a comparable BERTScore to other non-English languages, suggesting that its seemingly superior performance in traditional metrics reflects tokenization effects rather than genuine semantic alignment. In contrast, Chinese scores similarly to other non-English languages in metrics such as F1, BLEU, ROUGE-1, ROUGE-2, and ROUGE-L, yet records the lowest BERTScore at just 0.6086. This discrepancy indicates significant differences in structural and semantic expression between Chinese and other languages. Unlike many other languages, Chinese has less rigid grammatical requirements, allowing sentences with entirely different meanings to use similar words and word orders.
@@ 6.3.1 Mixed Training with English
Previous research often highlights that mixing English as a source and target language during training provides significant benefits for the target language due to English’s dominance in pretraining corpora and its rich semantic structure. However, our experiments on XMP challenge this assumption. For the LLaMA-3.1 8B model, mixing English with any other language consistently leads to negative performance impacts. Interestingly, for the LLaMA-3.1 70B model, this detrimental effect is mitigated, particularly when English is mixed with French. However, even in cases where certain languages exhibit improvements in specific metrics, these gains remain minimal, typically less than 0.5%. This suggests that while larger model capacities can temper the negative impacts of mixed training, the benefits derived from cross-lingual interaction are marginal and insufficient to fully leverage the potential of multilingual complementary ability in complex dialogue scenarios.
@@ 6.3.2 Cross-Language Benefits Are Asymmetrical
The experimental results for 70B-model reveal an intriguing asymmetry: while English typically fails to enhance the performance of other languages, other languages, such as French or German, often provide improvements when mixed with English. It might stem from the model’s reliance on English during pretraining, making it more adept at leveraging additional multilingual context for English-centric tasks while struggling to adapt English’s resources to enhance other languages effectively.
@@ 6.3.3 Effect of Linguistic Similarity
The relationship between linguistic similarity and cross-lingual gains appears to be more complex than previously assumed. Traditionally, languages within the same linguistic family, such as English and German, or French and English, are expected to exhibit significant cross-lingual transfer benefits when trained together. However, our experimental findings suggest that these linguistic connections do not guarantee consistent or significant cross-lingual improvements. For instance, the mixing of English and German often results in negative impacts on both languages’ performance, even though they are closely related in terms of linguistic roots. Conversely, the pairing of French with either German or English generally produces positive outcomes, albeit with varying degrees of improvement depending on the specific metrics and target languages.
The nuanced relationship highlights the need for deeper investigations into how linguistic similarity interacts with other variables in multilingual training, particularly in complex and dynamic tasks. It underscores that shared linguistic heritage, while beneficial, is not a deterministic factor for cross-lingual success, and the benefits of multilingual training are far from uniform across different language pairs
@@ 7 Conclusion
In this study, we introduced XMP, a high-quality multilingual parallel multi-party dialogue dataset, designed to address the gaps in existing datasets for complex dialogue scenarios. XMP provides extensive coverage of six languages to address the lack of comprehensive resources for complex multilingual dialogue generation. Using XMP as a benchmark, we evaluated the multilingual complementary ability of LLaMA-3.1 8B and 70B models, uncovering unexpected findings that challenge prior research. Mixing English with other languages during training often resulted in performance degradation, particularly for smaller models, while larger models like the 70B variant demonstrated better resilience and achieved marginal gains in specific cases, such as English-French mixing. Interestingly, linguistic similarity did not always yield mutual reinforcement; for instance, English and German showed negative impacts when mixed, whereas French consistently enhanced performance when paired with other languages. These insights underscore the limitations of current multilingual models and highlight the need for further research into linguistic properties and cross-lingual transfer, with XMP serving as a robust benchmark for advancing multilingual dialogue generation.
@@ Appendix A Preliminary Experiments
Before constructing the full dataset, we first conducted a small-scale collection of approximately 1,000 English multi-party dialogue samples. Each sample contained at least three speakers and covered a variety of topics to ensure diversity and complexity in the dialogues. To evaluate the large model’s multi-party dialogue response generation capability, we used In-Context Learning (ICL) testing. We found that as the number of examples in the ICL increased, there was virtually no improvement in the quality of the model’s generated responses. This was particularly evident with the English data, which led us to question how the large model would perform on relatively sparse, under-pretrained corpora. To investigate this further, we expanded the 1,000 English samples to include French, German, Italian, and Chinese.
@@
Following the approach of CCD et al., we applied different templates combined with ICL for further validation. Specially, we designed three different experimental methods to validate the model’s performance variations:
@@ B.1 Data Collection
LLaMA3 {CJK}UTF8gbsn一位毫无经验的制片人和他全新演员阵容准备创造电视史,可电视台的高层却紧盯着他们。几乎所有可能出错的地方都出了问题,真是糟透了。我是艾莎·哈里斯。
Human {CJK}UTF8gbsn一位未尝试的制片人带领一群毫无名气的演员,准备改写电视规则,可电视台高层却紧盯着他们的一举一动。结果几乎出错的地方都有错,真是糟糕了。我是艾莎·哈里斯。
@@ B.3 Data Segmentation
Appendix C Chosen of Translate API
@@
To further illustrate our reasoning for choosing the Google Translate API, the Table 7 and Table C present an actual translation snippet of a dialogue. In terms of word-level evaluation, all the translation APIs demonstrated accurate spelling without errors. Google’s translation of "cast of complete unknowns" as "{CJK}UTF8gbsn 一众无名小卒" is both vivid and precise. In contrast, Baidu’s "{CJK}UTF8gbsn网络服装" is a literal translation of "network suits," failing to capture the intended meaning. Similarly, DeepL’s "{CJK}UTF8gbsn西装革履" and iFlytek’s "{CJK}UTF8gbsn西装" translations of "network suits" are overly literal, lacking the semantic nuance of referring to "network executives."
@@ Appendix D Dataset Quality: Detailed evaluation criteria and the scoring results
To assess the quality of the dataset, we employed expert evaluation. Specifically, we randomly selected 500 conversations, covering various topics. The review was conducted based on the following aspects, with scores ranging from 1 to 5 for each metric. The final scores and detailed scoring criteria are shown in Table 9.
Metric Score Range Description Final Score
Conversation Diversity 1 Very Low Diversity: Narrow range of topics, repetitive dialogue structures, limited interaction patterns. 4.74
2 Low Diversity: Few different topics, minimal variation in interaction styles, some repetition.
3 Moderate Diversity: A reasonable range of topics, some diversity in interaction styles and contexts.
4 High Diversity: Diverse topics and contexts, multiple types of dialogues and varied interactions.
5 Very High Diversity: Rich variety of topics, multiple communication styles, diverse contexts and subjects.
Conversation Fluency 1 Very Low Fluency: Disjointed conversations with abrupt topic changes, frequent grammatical errors. 4.21
2 Low Fluency: Awkward transitions, slight disruptions in flow, minor grammatical issues.
3 Moderate Fluency: Clear dialogue with occasional awkward phrasing, understandable overall.
4 High Fluency: Natural flow, clear transitions between topics, minor errors not affecting overall fluency.
5 Very High Fluency: Seamless conversation with smooth transitions, grammatically correct throughout.
Annotation Quality 1 Very Low Quality: Responses are either too short or excessively long, often irrelevant or nonsensical. Responses frequently deviate from the previous context, making them incoherent or disjointed. 4.43
2 Low Quality: Responses are occasionally too short or too long, with some parts being irrelevant. There are noticeable inconsistencies with the previous context, leading to slight confusion in the conversation flow.
3 Moderate Quality: Responses are generally of appropriate length but may occasionally lack relevance or slightly stray from the context. They remain understandable and contextually coherent, with some minor lapses.
4 High Quality: Responses are of reasonable length, relevant, and mostly meaningful. There is good consistency with the previous dialogue, but small deviations or slight misunderstandings may still occur.
5 Excellent Quality: Responses are always appropriately lengthened, coherent, and directly related to the previous dialogue. They are contextually meaningful, contributing effectively to the conversation without unnecessary repetition or confusion.
Table 9: Detailed scoring criteria for dataset quality evaluation.

### v2

@@ 5.1 Dataset Split and Baseline Models
To benchmark the performance of state-of-the-art large language models (LLMs) on the XMP dataset, we selected three open-source instruction-tuned models as baselines: LLaMA 3.1 8B Instruct, Qwen 2.5 7B Instruct, and LLaMA 3.1 70B Instruct. These models were chosen for their demonstrated capabilities in multilingual dialogue response task. Their inclusion allows us to examine the trade-offs between model size, and response quality across six languages.
@@ 5.2 Task Setup
The generation task focuses on producing high-quality responses in complex multi-party dialogue scenarios.
Specially, Let $D={(H,S,R)}^{n}_{i=1}$ represents the multilingual dialogue dataset for complex scenarios, where $H$ is the dialogue history composed of multiple “Speaker: Utterance” pairs concatenated sequentially, $S$ is the current speaker, and $R$ is the ground truth response. To guide the model in generating speaker-specific responses, a prompt is dynamically constructed based on the current speaker, such as “What will Speaker-x say next?”. The dialogue history is then concatenated with this prompt to form the input for the model. During training, the model predicts the response $\hat{R}$, and the cross-entropy loss is computed between and the ground truth $L=-\frac{1}{m}{\textstyle\sum_{j=1}^{m}}R_{j}log\hat{R}_{j}$, where $m$ is the number of tokens in $R$.
@@ 5.3 Implementation Details
We utilized LLaMAFactory as the training framework and applied LoRA (Low-Rank Adaptation) Hu et al. (2021) for efficient fine-tuning. The training was conducted on NVIDIA A100 GPUs using the AdamW optimizer with a learning rate of $1e^{-5}$, a warm-up of 500 steps, and a batch size of 8. To enhance generalization, we standardized key entities in the dialogue history and replaced real entity data with domain-specific slot placeholders. For inference, we used vLLM, an optimized framework that ensures fast and scalable generation. Detailed training and inference commands, along with parameter settings, are provided in Appendix E.
@@ 6.1 Model Performance
The performance gain of LLaMA-3.1 70B over smaller models (e.g., 8B or Qwen-2.5 7B) is marginal, with improvements typically within 1% across most metrics (R2). This suggests that the emergent abilities often associated with larger models are not prominently reflected in this task. Instead, the slight advantage may stem from the larger models’ capacity to generate grammatically and lexically refined text, rather than demonstrating a better understanding of the task.
6.2 Structural and Semantic divergence
Japanese and Chinese both achieve relatively high scores on surface-level metrics like BLEU and ROUGE, Japanese even outperforming all other languages besides English, and Chinese ranking among the top as well. This can be attributed to factors such as syntactic regularity, frequent use of fixed expressions and kana in Japanese, and strong n-gram overlaps that benefit token-based evaluation.
However, as shown in Table 4, BERTScore tells a different story: Japanese aligns semantically with other non-English languages, while Chinese records the lowest score (0.6086), indicating significant structural and semantic divergence that models struggle to capture.
@@ 6.3.1 Mixed Training with English
Previous researches Whitehouse et al. (2023); Li et al. (2024) often highlight that mixing English as a source and target language during training provides significant benefits for the target language due to English’s dominance in pretraining corpora and its rich semantic structure. However, our experiments on XMP challenge this assumption. For the LLaMA-3.1 8B model, mixing English with any other language consistently leads to negative performance impacts. Interestingly, for the LLaMA-3.1 70B model, this detrimental effect is mitigated, particularly when English is mixed with French. However, even in cases where certain languages exhibit improvements in specific metrics, these gains remain minimal, typically less than 0.5%. This suggests that while larger model capacities can temper the negative impacts of mixed training, the benefits derived from cross-lingual interaction are marginal and insufficient to fully leverage the potential of multilingual complementary ability in complex dialogue scenarios. (R3)
@@ 6.3.2 Cross-Language Benefits Are Asymmetrical
The experimental results for 70B-model reveal an intriguing asymmetry: while English typically fails to enhance the performance of other languages, other languages, such as French or German, often provide improvements when mixed with English.
@@ 6.3.3 Effect of Linguistic Similarity
Both Eronen et al. (2023) and Aepli and Sennrich (2022) report that linguistic similarity plays a crucial role in cross-lingual transfer. Eronen et al. (2023) demonstrate this across sentiment analysis, and dependency parsing. Similarly, Aepli and Sennrich (2022) observe that in POS tagging, transfer between closely related language varieties yields significantly better performance than between unrelated pairs.
However, our experimental findings suggest that these linguistic connections do not guarantee consistent or significant cross-lingual improvements. For instance, the mixing of English and German often results in negative impacts on both languages’ performance, even though they are closely related in terms of linguistic roots. Conversely, the pairing of French with either German or English generally produces positive outcomes, albeit with varying degrees of improvement depending on the specific metrics and target languages.
The nuanced relationship highlights the need for deeper investigations into how linguistic similarity interacts with other variables in multilingual training, particularly in complex and dynamic tasks. It underscores that shared linguistic heritage, while beneficial, is not a deterministic factor for cross-lingual success, and the benefits of multilingual training are far from uniform across different language pairs.
6.4 Probing Context Sensitivity
Figure 3: Case Study
To examine how effectively the model leverages extended dialogue history, we conducted a controlled evaluation by randomly removing 30% of the context while retaining the final two turns. The rationale behind this setup is that if the model truly integrates long-range contextual information, removing part of the earlier context should noticeably impact its performance.
We applied this evaluation to the LLaMA 3 8B model on both English and Chinese subsets. Results are presented in Table 6. Overall, the differences across both languages were minimal.
These findings suggest that the model does not effectively utilize broader dialogue history. Its performance appears to rely primarily on recent turns, reflecting a limited capacity to model longer-range conversational dependencies.
Language Context F1 B-1 B-2 R-L
Chinese Full 18.37 11.60 4.68 5.44
Chinese - 30% 18.52 12.19 4.77 5.85
English Full 19.92 12.80 5.73 5.54
English - 30% 19.36 12.60 5.41 4.91
Table 6: Performance comparison with and without 30% context removal on the LLaMA 3 8B model. Context removal excludes the final two turns.
6.5 Case Study
To more intuitively demonstrate the current limitations of large language models in understanding multi-party conversations, we conduct a case study and analyze the model’s predictions. We found that the model’s responses often: 1. Misinterpreted conversational context and overlooked the roles of specific participants. 2. Lacked concrete factual content
We presented several cases in the Figure 3 . Due to character limitations, all examples shown here are in English — the model’s performance tends to degrade further in other languages.
In the first case, the model ignored the speaker Harris and abruptly shifted to a discussion of Biden’s plan. In contrast, the ground truth continued with a discussion of Harris’ media strategy, which was contextually appropriate. We attribute this error to amplified hallucination under complex dialogue settings.
In the second example, the model predicted a response like “I hope so,” which contributed nothing meaningful to the conversation, whereas the ground truth was a natural continuation of the ongoing topic. We believe this failure is likely due to: 1. the model struggles to maintain awareness of the broader conversational flow and context, particularly in multi-party settings; 2. current large models’ pretraining corpora still lack rich, multi-turn, real-world conversational data.
@@ 7 Conclusion
This study aims to explore the generalization capabilities of MLLM in complex dialogue scenarios. To enable this investigation, we construct XMP, a high-quality, parallel multilingual multi-party dialogue dataset. XMP serves as a critical resource to support rigorous evaluation in this challenging setting.
Through extensive experiments, we find: R1: MLLMs exhibit significant limitations in generalizing to multi-party dialogue. R2: Fine-tuning on XMP yields only marginal improvements, with larger models showing limited gains. R3: Contrary to prior studies, multilingual complementarity effects are inconsistent or even detrimental.
We hope our work and the XMP dataset will inspire further efforts toward building more context-aware, robust, and multilingual conversational AI.
@@ Appendix A Preliminary Experiments
Before constructing the full dataset, we first conducted a small-scale collection of approximately 1,000 English multi-party dialogue samples. Each sample contained at least three speakers and covered a variety of topics to ensure diversity and complexity in the dialogues. To evaluate the large model’s multi-party dialogue response generation capability, we used In-Context Learning (ICL) testing. We found that as the number of examples in the ICL increased, there was virtually no improvement in the quality of the model’s generated responses. This was particularly evident with the English data, which led us to question how the large model would perform on relatively sparse, under-pretrained corpora. To investigate this further, we expanded the 1,000 English samples to include French, German, Italian, and Chinese. The results are shown in Figure 5.
@@
Following the approach of Choenni et al. (2023) et al., we applied different templates combined with ICL for further validation. Specially, we designed three different experimental methods to validate the model’s performance variations:
@@ B.1 Data Collection
LLaMA3 {CJK}UTF8gbsn一位毫无经验的制片人和他全新演员阵容准备创造电视史,可电视台的高层却紧盯着他们。几乎所有可能出错的地方都会发生,真是糟透了。我是艾莎·哈里斯。
Human {CJK}UTF8gbsn一个毫无经验的制片人带领一群毫无名气的演员,准备改写电影史,可电视台高层却死死盯着他们的一举一动。结果几乎所有地方都出差错了,真是糟透了。我是艾莎·哈里斯。
@@ B.3 Data Segmentation
Appendix C Comparison of Translate Method
C.1 Human Evaluation for Different Translation Methods
Translation Method W-Level U-Level C-Level
Google Translate 92.6 86.9 84.5
Baidu Translate 88.2 73.1 74.7
DeepL Translate 88.7 70.8 73.4
iFLYTEK Translate 86.2 70.2 72.3
LLaMA 3.1 90.4 80.3 82.8
ChatGPT-4o 94.5 85.9 87.8
Human Translation 99.2 98.5 95.4
Table 8: Evaluation Results of Translation Methods. W-Level, U-Level and C-Level refer to Word-Level, Utterance-Level and Conversation-Level.
Language Method F1 B-1 B-2 R-L
Chinese GPT-4o 17.45 10.93 4.59 15.01
Chinese Google 17.61 11.01 4.55 14.97
German GPT-4o 16.97 11.60 4.91 13.88
German Google 17.12 11.51 4.95 13.84
Table 9: Performance comparison of LLaMA 3.1 trained on translations generated by GPT-4o vs. Google Translate (5k samples each).
For human evaluation, we began by selecting 50 representative samples that covered a diverse range of topics (e.g., law, politics, daily life), dialogue turns (ranging from 7 to 15 turns), and multi-party podcast segments. Each method translated the English transcripts into five other languages (Chinese, Japanese, French, German, and Italian), resulting in 300 translations. To ensure objectivity and reliability, three linguistics experts reviewed each translation in multiple rounds. The review process identified issues such as lexical misusage, semantic distortion, or contextual inconsistency in the translated outputs. Ultimately, the evaluation focused on three dimensions:
  • Conversation-level evaluation ensured contextual consistency and alignment with the original dataset.
As shown in Table 8, among Google Translate, DeepL, Baidu, iFLYTEK, LLaMA 3, ChatGPt-4o, both ChatGPT-4 and Google Translate have shown impressive results, with ChatGPT-4o slightly outperforming Google Translate. However, ChatGPT-4o requires careful handling of templates, as it may provide natural translations for some examples but sometimes offers literal translations that overlook specific contexts. Considering translation speed and cost-effectiveness, we opted for Google Translate as the more practical choice for large-scale multilingual tasks, just like prior works Lai et al. (2024); Liu et al. (2023).
C.2 Automatic Evaluation for Different Translation Methods
We further validated translation quality through a small-scale generation experiment, as shown in Table 9. A subset of approximately 5,000 English samples was translated into Chinese and German using both GPT-4o and Google Translate. We then fine-tuned LLaMA 3.1 using each translated dataset and compared dialogue generation performance. Results showed minimal performance differences between the two translation sources. These findings support the use of Google Translate as a practical and reliable solution for large-scale multilingual dialogue construction.
C.3 Case comparisons
Specific case comparisons of the translation tools are detailed below.
@@ C.3 Case comparisons
To further illustrate our reasoning for choosing the Google Translate API, the Table 7 and Table C.3 present an actual translation snippet of a dialogue. In terms of word-level evaluation, all the translation APIs demonstrated accurate spelling without errors. Google’s translation of "cast of complete unknowns" as "{CJK}UTF8gbsn 一众无名小卒" is both vivid and precise. In contrast, Baidu’s "{CJK}UTF8gbsn网络服装" is a literal translation of "network suits," failing to capture the intended meaning. Similarly, DeepL’s "{CJK}UTF8gbsn西装革履" and iFlytek’s "{CJK}UTF8gbsn西装" translations of "network suits" are overly literal, lacking the semantic nuance of referring to "network executives."
@@ Appendix D Dataset Quality: Detailed evaluation criteria and the scoring results
To assess the quality of the dataset, we employed expert evaluation. Specifically, we randomly selected 1000 conversations (200 samples for each language), covering various topics. The review was conducted based on the following aspects, with scores ranging from 1 to 5 for each metric.
Cross-lingual Consistency Analysis
use the Scoring Criteria below to assess how well the target-language text preserves the semantic meaning and conceptual structure of the English source text.
  • Semantic Equivalence: Accuracy in capturing key information, terminology, and meaning from the source.
  • Conceptual Alignment: Faithfulness to the overall theme, reasoning chain, and logical flow of the source text.
The Scoring Criteria and results are shown in Table 11 and Table 12.
Linguistic and Cultural Consistency
focuses on the quality of the target-language text in its own linguistic and cultural context.
  • Naturalness: Whether sentences are fluent, grammatically sound, and idiomatic.
  • Cultural Appropriateness: Alignment with cultural norms, conventions, and stylistic preferences.
The Scoring Criteria and result are shown in Table 13 and Table 14.
Score Semantic Equivalence Conceptual Alignment
1 Many key details lost or mistranslated; Significant deviation from source meaning Main topic is unclear or completely misaligned; Severe logical gaps or missing core concepts
2 Substantial inaccuracies in important information; Inadequate handling of key terms Overall theme is partially discernible, but reasoning steps are shaky; Noticeable gaps in coherence
3 Primary content is mostly correct; Some inaccuracies in less crucial details Main idea is largely maintained; Minor logical jumps but generally comprehensible
4 Nearly all key information is accurately preserved; Only minor misinterpretations The text closely follows the source theme; Reasoning and structure are coherent, with few gaps
5 All core information perfectly conveyed; No noticeable semantic errors Fully aligned with source logic; Clear and consistent reasoning; Nearly flawless
Table 11: Semantic Equivalence and Conceptual Alignment Scoring Guidelines
Target Language Samples Semantic Equivalence Conceptual Alignment
ZH (Chinese) 200 4.13 4.02
JA (Japanese) 200 4.17 4.11
FR (French) 200 4.32 4.25
DE (German) 200 4.05 3.92
IT (Italian) 200 3.93 3.88
Table 12: Quality Scores by Language
Score Naturalness Cultural Appropriateness
1 Numerous grammatical or lexical errors. Difficult to read or parse Severely mismatched with target-language cultural norms. Potentially offensive or highly unnatural expressions
2 Limited fluency; awkward or clumsy phrasing; Noticeable "translationese" Certain mismatches in cultural context; Expressions that appear slightly off or inauthentic
3 Mostly coherent sentences; Minor issues with word choice or syntax Understandable but not fully adapted to local norms; Some cultural references may be vague or misplaced
4 Generally fluent and easy to follow; Few or negligible foreign constructs Largely compatible with the target culture; Stylistically appropriate in most contexts
5 Exceptionally natural, no apparent translation artifacts; Flawless grammar Perfect cultural fit; expressions and style feel native; No inaccuracies or incongruities in cultural references
Table 13: Scoring Criteria for Naturalness and Cultural Appropriateness
Target Language #Samples Naturalness Cultural Appropriateness
ZH (Chinese) 200 3.94 4.10
JA (Japanese) 200 4.03 4.12
FR (French) 200 4.25 4.37
DE (German) 200 3.88 3.96
IT (Italian) 200 3.72 3.85
Table 14: Naturalness and Cultural Appropriateness Scores by Language